# Standard modules of affine Hecke algebras

### v1

@@ 2.5 Standard and essentially square-integrable representations
For generic $s$ and $\gamma(s)$, the standard modules $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ are simple. By [EvensMirkovic, Thm. 5.5], Lemma 3 already implies that $K(s,N,\rho)\simeq K(\gamma(s),N,\rho)$ in this case.
@@
A smooth admissible representation $\omega$ of $G^{\vee}(F)$ is essentially square integrable if it is of the form $\omega\otimes\nu$ for an unramified character $\nu$ and square-integrable modulo centre representation $\omega$ of $G^{\vee}(F)$.
Recall that standard representation $K(s,N,\rho)$ is an irreducible essentially square-integrable representation, if and only if, there is no proper parabolic subgroup of $G$ with Levi subgroup containing $(s,u)$, where $u=\exp(N)$ [KLDeligneLanglands, Thm. 8.3 (a)].
@@
More generally, if $\pi$ is a standard $H$-module, so is $\gamma\cdot\pi$, and twist by $\gamma$ also preserves genericity of the Langlands quotients.
@@
Now, in general, $\gamma$ of course does not commute with complex conjugation. In the case at hand, though, square-integrability of $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ implies that $s_{1}$ and $\gamma(s_{1})$ belongs to the maximal compact subgroup of $T^{\vee}$. Therefore
@@
with $\lambda(\gamma(s_{1}))\in U(1)\subset\mathbb{C}^{\times}$ for any weight $\lambda$, and, by hypothesis, likewise for $s_{1}$ (note $\gamma(s_{1})=\gamma(s)\phi(\gamma(q^{-1/2},q^{1/2})=\gamma(s)\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2})$). Therefore
@@
3 Essentially discrete series representations of $\mathrm{GL}_{n}$
@@ 3.1 Types and Hecke algebras
In this section, we give our proof of Theorem 2 for $G=\mathrm{GL}_{n}$. We first recall
@@
Let $F$ be a non-archimedean local field and let $G=\mathrm{GL}_{n}(F)$; let $q$ be the cardinality of the residue field of $F$. For any Bernstein block $\mathbf{Rep}_{\mathfrak{s}}(G)$ of the category of smooth representations $\mathbf{Rep}(G)$ of $G$, there are integers $\left\{m_{i}\right\}_{i}$ and $\left\{f_{i}\right\}_{i}$ and an equivalence of categories
$\mathbf{Rep}_{\mathfrak{s}}(G)\simeq\bigotimes_{i}H_{m_{i}}(q^{f_{i}}),$
@@
Theorem 6 ([BHK], Thm. 4.3).
The equivalences of Theorem 5 preserve Plancherel measures up to a constant factor, and hence preserve discrete series and tempered representations.
Note that Theorem 6 applies to any type, not only the ones constructed in [BushKutzBook], [BushKutzComp].
As remarked in the introduction, if $\sigma$ is a supercuspidal representation of $G$, then clearly $\gamma\cdot\sigma$ is again supercuspidal. The compatibility with normalized parabolic induction, up to unramified twist of the induced representation, of $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}(-)$ [KSV, (5.12)] then implies that the auto-equivalence $\pi\mapsto\gamma\cdot\pi$ of $\mathbf{Rep}(G)$ sends the Bernstein block for an inertial class $\mathfrak{s}=[M,\sigma]_{G}$ to the Bernstein block for $\mathfrak{s}_{\gamma}=[M,\gamma\cdot\sigma]_{G}$. As remarked above, $T,\mathrm{triv}]_{G}=[T,\gamma\cdot\mathrm{triv}]_{G}$, but in general $\mathfrak{s}_{\gamma}$ labels a new block. If $(K,\rho)$ is a type for $\mathfrak{s}$, then $(K,\gamma\cdot\rho)$ is a type for $\mathfrak{s}_{\gamma}$, and we have an isomorphism of rings
@@
Rewriting this in terms of affine Hecke algebras, we have isomorphisms
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}H_{m_{i}}(q^{f_{i}})\overset{\sim}{\to}H(G,K,\gamma\cdot\rho)\simeq\bigotimes_{i}H_{m_{i}}(q^{f_{i}}),$
where the second isomorphism is induced by the auto-equivalence $\pi\mapsto\gamma\cdot\pi$.
Now suppose that $\pi$ is a simple essentially square-integrable representation in $\mathbf{Rep}(G)_{\mathfrak{s}}$. Then $\pi$ corresponds to a $H(G,K,\rho)$-module
@@
3.2 Proof of Theorem 2 for $G=\mathrm{GL}_{n}$
To show that $\gamma\cdot\pi$ is again essentially discrete series, it suffices by Section 3.1 to show that the $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$-module (13) is isomorphic via $\bigotimes_{i}\Phi(\gamma)$ to $\bigotimes_{i}K(\gamma(s_{i}),N_{i})$ under the second isomorphism above to $\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$. This also implies the statement about formal degrees, as formal degrees of essentially discrete series representations of affine Hecke algebras of type $\mathrm{GL}_{n}$ depend only on $N$. Therefore, it suffices to prove
@@ 3.2 Proof of Theorem 2 for $G=\mathrm{GL}_{n}$
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle s\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{(s,q)})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle\gamma(s)\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q)})$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{\gamma(s)})_{\mathbb{Q}},$$\scriptstyle{i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{i_{\gamma(s)}\circ\mathrm{Forg}_{\left\langle\gamma(s)\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{id}\otimes(-\otimes(1\boxtimes\lambda_{s}^{-1}))}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes(-\otimes(1\boxtimes\lambda_{\gamma}(s)^{-1}))}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}(v)}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}(\Phi(\gamma)v)}$$\scriptstyle{\Phi(\gamma)}$
@@
In the diagram, we write $\left\langle s\right\rangle$ for the smallest closed subgroup of $G$ containing $s$. Note also that that $H_{*}(\mathrm{St}^{s};\mathbb{Q})$ and $H_{*}(\mathrm{St}^{\gamma(s)};\mathbb{Q})$ are isomorphic to $A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$ and $A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$, respectively. This follows from [Fulton, Cor. 18.3.2] and [CG, Thm. 6.2.4] (the proof in loc. cit. works for any field coefficients). The Künneth formula for homology and [CG, (5.2.4)] reduce the general case to products of the case of the Proposition.
@@
It is enough to check commutativity of the square on elements of the form $1\otimes 1\otimes\mathcal{F}$ for generators $\mathcal{F}$ of $K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})$. The bundles $\mathcal{O}_{\lambda}$ on the diagonal component $\Delta\tilde{\mathcal{N}}\subset\mathrm{St}$, pulled back from the line bundles $\mathcal{O}(\lambda)$ on $\mathcal{B}$ generate the Bernstein subalgebra. We checked commutativity for the Bernstein subalgebra in Section 2.4.
@@
    Taking a generator $\mathcal{L}_{\sigma}$, the character that appears in the $K(\mathrm{pt}/\left\langle s\right\rangle)$-coordinate upon applying $i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}$ is of the form $\alpha(\varphi_{B_{j}}(s))$ as in (9), where $\mathcal{B}_{j}$ is a connected component of $\mathcal{B}^{s}$ the proof of Lemma 3. We conclude again that the square commutes. Finally, note that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}_{\sigma}]=\mathcal{L}_{\sigma}$, as the Schubert varieties and their conormal bundles are defined over $\mathbb{Q}$. This holds as equivariant sheaves: concretely, the equivariant structure is a vector bundle morphism $G\times\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}\to\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}$ over $T^{*}_{\bar{Y}_{\sigma}}(\mathcal{B}\times\mathcal{B})$ commuting with $G$ action and fibrewise linear, and this morphism is defined over $\mathbb{Q}$. Therefore this square commutes.
@@
    where $N=\bigoplus_{\alpha\in\mathrm{sp}(N)}N_{\alpha}$ is the $\left\langle s\right\rangle$-weight decomposition of the normal bundle $N$ to $\tilde{\mathcal{N}}^{s}$, and that, as remarked in loc. cit.,
    $\mathcal{O}_{\tilde{\mathcal{N}}}\boxtimes\lambda_{s}^{-1}$
    exists in complexified $K$-theory [CG, Cor. 5.11.3]. It is then clear that $\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}\lambda_{s}=\lambda_{\gamma(s)}$, and hence the same is true for the inverse classes. Here, we again use that the considerations of (9) apply also the bundles $\mathcal{L}_{\sigma}$, by their construction recalled above. Therefore this square commutes.
@@
    where $p_{i}$ are the respective projections $\mathrm{St}^{s}\to\tilde{\mathcal{N}}^{s}$ and we view $v$ as a cycle in the (smooth, by [CG, Lem. 5.11.1] variety $\tilde{\mathcal{N}}^{s}$ (to which Theorem 6.2.4 of op. cit. also applies). Now we apply base-change [Fulton, Prop. 1.7] (see [CG, §2.7] for the corresponding statements for homology throughout) to the diagram

### v2

@@ 2.5 Standard and essentially square-integrable representations
For generic $s$ and $\gamma(s)$, the standard modules $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ are simple. For $G$ simply connected, by [EvensMirkovic, Thm. 5.5], Lemma 3 already implies that $K(s,N,\rho)\simeq K(\gamma(s),N,\rho)$ in this case.
@@
A smooth admissible representation $G^{\vee}(F)$ is essentially square integrable if it is of the form $\omega\otimes\nu$ for an unramified character $\nu$ and square-integrable modulo centre representation $\omega$ of $G^{\vee}(F)$.
Recall that a standard representation $K(s,N,\rho)$ is an irreducible essentially square-integrable representation, if and only if, there is no proper parabolic subgroup of $G$ with Levi subgroup containing $(s,u)$, where $u=\exp(N)$ [KLDeligneLanglands, Thm. 8.3 (a)]. As recalled in [Reeder, §2.1], this is equivalent to $Z_{G}(s)$ being semisimple with $N$ a distinguished nilpotent in $\operatorname{\mathrm{Lie}}(Z_{G}(s))$. Hence the eigenvalues of $s$ are roots of unity by Kac’s classification.
@@
More generally, if $\pi$ is a standard $H$-module, so is $\gamma\cdot\pi$, and twist by $\gamma$ also preserves genericity of the Langlands quotients (but with respect to the Whittaker datum for the $\gamma$-twisted generic character.
@@
Now, in general, $\gamma$ of course does not commute with complex conjugation. In the case at hand, though, square-integrability of $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ implies that $s_{1}$ and $\gamma(s_{1})$ belong to the maximal compact subgroup of $T^{\vee}$. Therefore
@@
with $\lambda(\gamma(s_{1}))\in U(1)\subset\mathbb{C}^{\times}$ for any weight $\lambda$, and, by hypothesis, likewise for $s_{1}$, noting that
$\gamma(s_{1})=\gamma(s)\phi(\gamma(q^{-1/2},q^{1/2})=\gamma(s)\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2}).$
Therefore
@@
Corollary 2.
If $L_{s,N,\rho}$ is the simple $H_{I}$-module associated to $(s,N,\rho)$ by [KLDeligneLanglands] or [CG], then $\gamma\cdot L_{s,N,\rho}=L_{\gamma(s),N,\rho}$.
Proof.
By definition, $L_{s,N,\rho}$ is the unique simple quotient $H$-module of $K(s,N,\rho)$. By exactness of $\gamma\cdot(-)$, $\gamma\cdot K(s,N,\rho)$ has unique simple quotient $\gamma\cdot L(s,N,\rho)$. By Corollary 1, $\gamma\cdot K(s,N,\rho)=K(\gamma(s),N,\rho)$. ∎
3 Essentially discrete series representations of inner forms of $\mathrm{GL}_{n}$
@@ 3.1 Types and Hecke algebras
In Section 3.1 only, we reverse our conventions on $G$ vs $G^{\vee}$.
In this section, we give our proof of Theorem 2 for $G$ an inner form of $\mathrm{GL}_{n}$. We first recall
@@
Let $F$ be a non-archimedean local field and let $G=\mathrm{GL}_{n}(F)$; let $q$ be the cardinality of the residue field of $F$. Then
  1. 1.
    For any Bernstein block $\mathbf{Rep}_{\mathfrak{s}}(G)$ of the category of smooth representations $\mathbf{Rep}(G)$ of $G$, there are integers $\left\{m_{i}\right\}_{i}$ and $\left\{f_{i}\right\}_{i}$ and an equivalence of categories
    $\mathbf{Rep}_{\mathfrak{s}}(G)\simeq\bigotimes_{i}H_{m_{i}}(q^{f_{i}})-\mathbf{Mod},$ (12)
@@
  2. 2.
    If $\mathbf{Rep}_{\mathfrak{s}}(G)$ contains an essentially discrete series representation, then on the right hand side of (12) appears a single Hecke algebra $H_{m}(q^{f})$, and the equivalence (12) induces a bijection of essentially discrete series representations preserving formal degrees up to explicit scalar factors.
The statements of the second part are Corollary 8.5.11, Theorem 7.7.1, and Corollary 7.7.11 of [BushKutzBook], respectively.
Now let $G$ be an inner form of $\mathrm{GL}_{n}$, so that $G(F)=\mathrm{GL}_{n}(D)$ for a division algebra $D$ with centre $F$.
Theorem 6 ([Secherre], [SecherreStevens]).
Let $G$ be an inner form of $\mathrm{GL}_{n}$ and $L$ be a Levi subgroup, so that $L(F)=\prod_{i}\mathrm{GL}_{n_{i}}(D)^{e_{i}}$ with $\sum_{i}n_{i}e_{i}=n$ and let $\omega=\bigotimes_{i}\omega_{i}^{\otimes e_{i}}$ be a supercuspidal representation of $L(F)$ for $\omega_{i}$ a supercuspidal representation of $\mathrm{GL}_{n_{i}}(D)$, grouped such that $\omega_{i}$ is not inertially equivalent with $\omega_{j}$ if $i\neq j$. Let $\mathfrak{s}=[L,\omega]$ be the corresponding inertial equivalence class and $\mathbf{Rep}(G)_{\mathfrak{s}}$ the resulting Bernstein block. Then
  1. 1.
    There is an equivalence of categories
    $\mathbf{Rep}(G)_{\mathfrak{s}}\simeq\bigotimes_{i}H_{e_{i}}(q^{f_{i}})-\mathbf{Mod}$ (13)
    for integers $f_{i}$, where $H_{e_{i}}(q^{f_{i}})$ is an affine Hecke algebra of type $\mathrm{GL}_{e_{i}}$ with parameter $q^{f_{i}}$, and the tensor product is taken over $\mathbb{C}$.
  2. 2.
    If $\mathbf{Rep}_{\mathfrak{s}}(G)$ contains an essentially discrete series representation, then $\mathfrak{s}$, then on the right hand side of (13) appears a single Hecke algebra $H_{e}(q^{f})$, and the equivalence (12) induces a bijection of essentially discrete series representations preserving formal degrees up to explicit scalar factors.
The fact that essentially square-integrable representations can only contain simple types follows from [Secherre], [SecherreStevensIV] (c.f. [SecherreV, p.3]). As noted by Solleveld [SolleveldCompletion], the remaining statements follow as in [BushKutzBook, §7] (c.f. the discussion at the start of Section 7.7 of loc. cit.), once one knows that the isomorphism of affine and typical Hecke algebras behind (13) preserves the *-involution and the trace, as with these two ingredients one builds the pairing after (7.7.3) of loc. cit.. This preservation is checked by Solleveld [SolleveldCompletion, Thm. 4.3].
As remarked in the introduction, if $\sigma$ is a supercuspidal representation of $G(F)$, then clearly $\gamma\cdot\sigma$ is again supercuspidal. The compatibility with normalized parabolic induction, up to unramified twist of the induced representation, of $\gamma\cdot(-)$ [KSV, (5.12)] then implies that the exact auto-equivalence $\pi\mapsto\gamma\cdot\pi$ of $\mathbf{Rep}(G)$ sends the Bernstein block for an inertial class $\mathfrak{s}=[L,\sigma]_{G}$ to the Bernstein block for $\mathfrak{s}_{\gamma}=[L,\gamma\cdot\sigma]_{G}$. As remarked above, $T,\mathrm{triv}]_{G}=[T,\gamma\cdot\mathrm{triv}]_{G}$, but in general $\mathfrak{s}_{\gamma}$ labels a new block. Note that $\mathfrak{s}$ is homogeneous if and only if $\gamma\cdot\mathfrak{s}$ is.
If $(K,\rho)$ is a type for $\mathfrak{s}$, then $(K,\gamma\cdot\rho)$ is a type for $\mathfrak{s}_{\gamma}$, and we have an isomorphism of rings
@@
Recall that the construction of types in [BushKutzBook], [BushKutzComp], [Secherre] [SecherreStevens] is done with respect to a chosen additive character $\psi_{F}\colon F\to\mathbb{C}^{\times}$, but that neither the notion of essentially square-integrable $G(F)$-representation, essentially-square-integrable module over an affine Hecke algebra, refer to this choice, and likewise with standard representations and standard modules.
Lemma 4.
The type $(K,\gamma\cdot\rho)$ is again a type obtained by the constructions of Theorem 5 and 6, but with respect to the additive character $\gamma\circ\psi$. (c.f. Corollary 1.)
Proof.
We adopt the terminology and notation of [BushKutzBook], [Secherre].
As in the proof of Theorem 7, it is enough to check this for simple types, corresponding to a single tensor factor in either of Theorems 5 or 6. Hence we consider simple strata $\mathfrak{A},n,0,\beta]$. The definition of simplicity refers to $\psi$, but the sets of simple strata are the same for $\psi$ and $\psi\circ\gamma$. Note also that the *- and $\dagger$-duals of Sections 1.1.4 and 1.3.6 of [BushKutzBook], respectively, are the same for $\psi$ and $\gamma\circ\psi$, and that Section 1.3 of op. cit. depends only on the (unchanged by $\gamma)$ conductor of $\psi$, and the coherences provided by Theorem 3.6.1 of op. cit. are $\mathrm{Aut}(\mathbb{C})$-equivariant.
Now, the representation $\rho$ is constructed as $\rho=\kappa\otimes\sigma_{0}$ is the tensor product of a homogeneous cuspidal representation of the finite group of Lie type obtained a quotient of a certain open compact subgroup $J=J(\beta,\mathfrak{A})$, and a character $\kappa$ obtained as follows: First, $\kappa$ is an extension to $J$ of a representation $\eta(\theta)$, in turn an extension of a simple character (in the sense of [BushKutzBook]) $\theta$ of the group $H^{1}(\beta,\mathfrak{A})$.
The characterization of $\kappa$ in terms of $\eta$ in terms of intertwining sets implies that $\gamma\cdot\kappa(\eta)=\kappa(\gamma\cdot\eta)$, and in turn the unicity of $\eta(\theta)$ implies $\gamma\cdot\eta(\theta)=\eta(\gamma\cdot\theta)$. Tracing the definition of special character back though [SecherreI] to [BushKutzBook, (3.2.1)], we see that if $\theta$ is a special character of $H^{m+1}(\beta)$ defined with respect to $\psi$, then $\gamma\cdot\theta$ is a simple character of the same subgroup for the same stratum, only with respect to $\gamma\circ\psi$.
Clearly cuspidality of $\sigma_{0}$ is preserved by $\gamma\cdot(-)$. ∎
Therefore to both $H(G,K,\rho)$ and $H(G,K,\gamma\cdot\rho)$ we can apply
Theorem 7 ([SolleveldCompletion], p. 43).
The equivalences of categories in Theorem 5 and 6 preserve standard representations.
We emphasize that this is a non-trivial condition because in general the root datum $\mathcal{R}$ associated to some $H(G,K,\rho)$ is not just that of $G$.
Now suppose that $\pi$ is a standard representation in $\mathbf{Rep}(G)_{\mathfrak{s}}$. By Theorem 7, $\pi$ corresponds to a $H(G,K,\rho)$-module
@@
3.2 Proof of Theorems 1 and 2 for inner forms of $\mathrm{GL}_{n}$
We return to our convention that $G^{\vee}$ is the $p$-adic group.
To show that $\gamma\cdot\pi$ is again standard, it suffices by Section 3.1 to show that the $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$-module (15) is isomorphic via $\bigotimes_{i}\Phi(\gamma)$ to $\bigotimes_{i}K(\gamma(s_{i}),N_{i})$ under the second isomorphism above to $\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$. Therefore, it suffices to prove
@@ 3.2 Proof of Theorems 1 and 2 for inner forms of $\mathrm{GL}_{n}$
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle s\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{(s,q)})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle\gamma(s)\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q)})$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{\gamma(s)})_{\mathbb{Q}},$$\scriptstyle{i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{i_{\gamma(s)}^{*}\circ\mathrm{Forg}_{\left\langle\gamma(s)\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{id}\otimes\left(-\otimes\left(1\boxtimes\lambda_{s}^{-1}\right)\right)}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\left(-\otimes\left(1\boxtimes\lambda_{\gamma(s)}^{-1}\right)\right)}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}(v)}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}\left(\Phi\left(\gamma\right)v\right)}$$\scriptstyle{\Phi(\gamma)}$
@@
In the diagram, we write $\left\langle s\right\rangle$ for the smallest closed subgroup of $G$ containing $s$. Note also that that $H_{*}(\mathrm{St}^{s};\mathbb{Q})$ and $H_{*}(\mathrm{St}^{\gamma(s)};\mathbb{Q})$ are isomorphic to $A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$ and $A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$, respectively. This follows from [Fulton, Cor. 18.3.2] and [CG, Thm. 6.2.4] (the proof in loc. cit. works for $\mathbb{Q}$-coefficients, by [CG, Rem. 8.1.10]). The Künneth formula for homology and [CG, (5.2.4)] reduce the general case to products of the case of the Proposition.
Of course, in the case of essentially square-integrable representations, only a single tensor factor appears, and Proposition 1 also implies the statement about formal degrees, as formal degrees of essentially discrete series representations of affine Hecke algebras of type $\mathrm{GL}_{r}$ depend only on $N$, and all of the scalar factors in [BushKutzBook, Cor. 7.7.11] are $\mathrm{Aut}(\mathbb{C})$-fixed.
@@
It is enough to check commutativity of the square on elements of the form $1\otimes 1\otimes\mathcal{F}$ for generators $\mathcal{F}$ of $K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})$. The bundles $\mathcal{O}_{\lambda}$ on the diagonal component $\Delta\tilde{\mathcal{N}}\subset\mathrm{St}$, pulled back from the line bundles $\mathcal{O}(\lambda)$ on $\mathcal{B}$, generate the Bernstein subalgebra. We checked commutativity for the Bernstein subalgebra in Section 2.4.
@@
    Taking a generator $\mathcal{L}_{\sigma}$, the character that appears in the $K(\mathrm{pt}/\left\langle s\right\rangle)$-coordinate upon applying $i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}$ is of the form $\alpha(\varphi_{B_{j}}(s))$ as in (9), where $\mathcal{B}_{j}$ is a connected component of $\mathcal{B}^{s}$ as in the proof of Lemma 3. We conclude again that the square commutes. Finally, note that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}_{\sigma}]=\mathcal{L}_{\sigma}$, as the Schubert varieties and their conormal bundles are defined over $\mathbb{Q}$. This holds as equivariant sheaves: concretely, the equivariant structure is a vector bundle morphism $G\times\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}\to\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}$ over $T^{*}_{\bar{Y}_{\sigma}}(\mathcal{B}\times\mathcal{B})$ commuting with $G$ action and fibrewise linear, and this morphism is defined over $\mathbb{Q}$. Therefore this square commutes.
@@
    where $N=\bigoplus_{\alpha\in\mathrm{sp}(N)}N_{\alpha}$ is the $\left\langle s\right\rangle$-weight decomposition of the normal bundle $N$ to $\tilde{\mathcal{N}}^{s}$, and that $\mathcal{O}_{\tilde{\mathcal{N}}}\boxtimes\lambda_{s}^{-1}$ exists in complexified $K$-theory [CG, Cor. 5.11.3]. It is then clear that $\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}\lambda_{s}=\lambda_{\gamma(s)}$, and hence the same is true for the inverse classes. Here, we again use that the considerations of (9) apply also the bundles $\mathcal{L}_{\sigma}$, by their construction recalled above. Therefore this square commutes.
@@
    where $p_{i}$ are the respective projections $\mathrm{St}^{s}\to\tilde{\mathcal{N}}^{s}$ and we view $v$ as a cycle in the smooth, by [CG, Lem. 5.11.1], variety $\tilde{\mathcal{N}}^{s}$ (to which Theorem 6.2.4 of op. cit. also applies). Now we apply base-change [Fulton, Prop. 1.7] (see [CG, §2.7] for the corresponding statements for homology throughout) to the diagram